OpenStudy (anonymous):

2Ix+2I-1=5 Which represents the solution t the absolute value equation? a.x=4 or x=8 b. x=1 or x=-1 c. x=1 or x=-5 d. x=5 or x=-5

OpenStudy (anonymous):

start by isolating the variable we're solving for. in this case, we have some extra things attached to the variable. \(2|x+2|-1=5\) step 1: add 1 to both sides \(2|x+2|=6\) step 2: divide both sides by 2 \(|x+2|=3\) then we can set up 2 equations: \(x+2=3\\x+2=-3\) solve: \(x= 1,-5\)